The sum of three consecutive odd integers is 75. Find the integers.
Answer by math_helper(2461)


The middle number is 75/3 = 25, thus 23,25, and 27 are the three consecutive odd integers.
---
The workout: n-2, n, n+2 are the three consecutive odd integers
3n = 75
n = 25 --> n-2 = 23, and n+2 = 27